A simple random sample of 20 items resulted in a sample mean of 15. The population standard deviation is σ = 6. Round your answers to two decimal places a. what is the standard error of the mean, σ ? At 95% confidence, what is the margin of error?
Answer #1

Solution :

Given that,

\bar x = 15

\sigma = 6

n = 20

A ) Standard error   of the mean is \sigma \bar x

=  (\sigma\bar x/\sqrtn)

= 6 \sqrt 20

= 1.34

Standard error = 1.34

B ) At 95% confidence level the z is ,

\alpha  = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.960

Margin of error = E = Z\alpha/2* (\sigma\bar x\sqrt/n)

= 1.960 * (6 / \sqrt 20 )

= 2.63

Margin of error = E = 2.63